### v1

# Surgical Gaussian Surfels: Highly Accurate Real-time Surgical Scene Rendering

Abstract
Accurate geometric reconstruction of deformable tissues in monocular endoscopic video remains a fundamental challenge in robot-assisted minimally invasive surgery. Although recent volumetric and point primitive methods based on neural radiance fields (NeRF) and 3D Gaussian primitives have efficiently rendered surgical scenes, they still struggle with handling artifact-free tool occlusions and preserving fine anatomical details. These limitations stem from unrestricted Gaussian scaling and insufficient surface alignment constraints during reconstruction. To address these issues, we introduce Surgical Gaussian Surfels (SGS), which transforms anisotropic point primitives into surface-aligned elliptical splats by constraining the scale component of the Gaussian covariance matrix along the view-aligned axis. We predict accurate surfel motion fields using a lightweight Multi-Layer Perceptron (MLP) coupled with locality constraints to handle complex tissue deformations. We use homodirectional view-space positional gradients to capture fine image details by splitting Gaussian Surfels in over-reconstructed regions. In addition, we define surface normals as the direction of the steepest density change within each Gaussian surfel primitive, enabling accurate normal estimation without requiring monocular normal priors. We evaluate our method on two in-vivo surgical datasets, where it outperforms current state-of-the-art methods in surface geometry, normal map quality, and rendering efficiency, while remaining competitive in real-time rendering performance. We make our code available at https://github.com/aloma85/SurgicalGaussianSurfels
Keywords:
Gaussian Splatting 3D Surface Reconstruction Minimally Invasive Surgery.
1 Introduction
Reconstructing dynamic surgical scenes from endoscopic videos is a crucial, yet complex task in robot-aided minimally invasive surgery. Accurate reconstructions improve the precision of surgical instrument manipulation and serve as a foundation for clinical applications such as AR/VR surgical simulation, medical training, and complete medical automation [2]. Factors such as deformable tissues, limited viewing angles, surgical tool occlusion, and severe specularity currently thwart existing 3D surface rendering methods [4].
SGS Result Reference SGS Result Reference
Figure 1: Qualitative results of our method SGS on the StereoMIS [5] Intestine and Liver dataset. We remove surgical tools and render artifact-free surgical tissue.
Advancements in 3D Gaussian Splatting (3DGS) [8] have shown promise in rendering and reconstructing 3D surgical scenes efficiently. Despite these advancements, 3DGS struggles with accurate geometric reconstructions, primarily due to the non-zero thickness of Gaussian point primitives and the ambiguity in normal directions due to their ellipsoid-like shape [1]. To address these limitations, Gaussian Surfels have been proposed, which flatten 3D Gaussian points into 2D ellipses by setting the z-scale to zero [1].
In this paper, we introduce Surgical Gaussian Surfels (SGS), a novel method that combines the surface alignment properties of Gaussian Surfels with the dynamic modeling capabilities of Gaussian Splatting to reconstruct deformable tissues in endoscopic videos. Occluding surgical instruments introduce challenges in reconstruction by obstructing the view of underlying tissues and introducing motion-appearance ambiguities in single-viewpoint scenarios. We address this by incorporating depth priors and a binary motion mask during Gaussian Surfel initialization and training, effectively filtering out occluding instruments and improving reconstruction accuracy.
Our contributions are as follows: (1) We propose SGS, a point-primitive-based rendering technique that takes advantage of the surface-alignment benefits and the high-quality rendering abilities of Gaussian Surfels to achieve high-fidelity reconstruction of dynamic surgical scenes; (2) By defining surface normals as the direction of the steepest density change within each surfel primitive, we accurately predict the orientation of surgical tissue without relying on monocular normal priors, thereby enhancing reconstruction geometry and mitigating shape-radiance ambiguities in highly specular areas; (3) We introduce a Projection-Based Iterative Mask Integration (PIMI) method to initialize point clouds in surgical scenes, resolving gaps and inaccuracies caused by incomplete depth data through confidence-weighted depth aggregation and refined color mapping; (4) SGS outperforms prior implicit and explicit rendering techniques in both quantitative and qualitative experiments.
2 Related Work
Recent advances in endoscopic 3D reconstruction have been driven by neural-radiance fields and explicit representations. EndoNerf [16] introduces a dual network architecture that combines implicit neural fields with mask-guided ray casting for tool occlusion and depth-cueing ray marching for single-viewpoint reconstruction. EndoSurf [20] advances this framework by representing the canonical scene using signed distance functions (SDF). Neural LerPlane [18] decomposes 4D surgical scenes by factorization into explicit 2D planes of static and dynamic fields to address the computational overhead of volumetric neural fields. EndoGaussian [10] marks a shift towards explicit representations by introducing Holistic Gaussian Initialization (HGI). Endo-4DGS [7] leverages the Depth-Anything foundation model for robust initialization and implements confidence-guided learning to handle uncertain depth measurements in deformable tissues. SurgicalGaussian [17] addresses the limitation of planar structures in encoding complex motion fields by using a deformation network that decouples motion and geometry. Our method builds on and surpasses these prior methods in artifact-free surgical tool removal and tissue reconstruction.
3 Method
Figure 2: Method pipeline. The full video sequence frames, Binary Mask, Depth Map is passed through PIMI; (a) 2D points are projected to 3D then the tissue surface is represented as Gaussian Surfels with scaling matrix $\mathbf{S}_{k}$, tangent vectors $\mathbf{t}_{u}$ and $\mathbf{t}_{v}$, and position $\mathbf{x}_{k}$; (b) Deformation prediction MLP takes Surfel position $\mathbf{x}_{k}$ and normalized timestamp $\mathbf{t}$ then predicts the change in position $\delta{x}$, scale $\delta{s}$, and rotation $\delta{q}$; (c) Optimization of Surgical Gaussian Surfels. Pipeline outputs an occlusion-free rendered surgical scene, normal map, and depth map.
Our pipeline, shown in Fig. 2, takes as input a complete monocular endoscopic video sequence, the corresponding binary occlusion masks, and depth maps. The time component $t_{i}$ of each frame $i$ is normalized to $t_{i}=i/T$. We initialize our Surgical Gaussian Surfels from the PIMI (Sec. 3.2) reconstructed point cloud, representing each point as a surface-aligned elliptical splat characterized by a constrained covariance matrix, opacity value, and view-dependent color features represented by spherical harmonic coefficients. To model tissue deformation, we employ a lightweight MLP that predicts per-surfel motion fields while maintaining temporal consistency. The network is set to minimize the photometric differences between the input frames and the rendered views. We introduce several regularization terms to enforce surface smoothness and geometric consistency in occluded regions (Sec. 3.3). Our pipeline enables real-time rendering of dynamic surgical scenes while preserving fine surgical tissue details.
3.1 Preliminaries and Surgical Gaussian Surfels
3D Gaussian Splatting represents a scene as a set of anisotropic 3D Gaussian primitives $\{\mathcal{G}_{k}\mid k=1,\dots,K\}$, each parameterized by its position $\mathbf{p}_{k}\in\mathbb{R}^{3}$, opacity $\alpha_{k}\in[0,1]$, and scaling matrix $\mathbf{S}_{k}\in\mathbb{R}^{3\times 3}$. The covariance matrix of $\mathcal{G}_{k}$ in the world space is defined as $\boldsymbol{\Sigma}_{k}=\mathbf{O}_{k}\operatorname{diag}(\mathbf{s}_{k})\mathbf{O}_{k}^{T}$, where $\mathbf{s}_{k}\in\mathbb{R}^{3}$ is a scaling vector and $\mathbf{O}_{k}\in\mathbb{R}^{3\times 3}$ is a rotation matrix parameterized by a quaternion. The Gaussian function is defined as $\mathcal{G}_{k}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{k})^{T}\boldsymbol{\Sigma}_{k}^{-1}(\mathbf{x}-\mathbf{p}_{k})}$. During rendering, Gaussians undergo a view transformation defined by rotation $\mathbf{R}$ and translation $\mathbf{t}$, where $\mathbf{p}_{k}^{\prime}=\mathbf{R}\mathbf{p}_{k}+\mathbf{t},\ \boldsymbol{\Sigma}_{k}^{\prime}=\mathbf{R}\boldsymbol{\Sigma}_{k}\mathbf{R}^{T}$ represent the view-space position and the covariance matrix, respectively. They are then projected to ray space through a local affine transformation $\boldsymbol{\Sigma}_{k}^{2D}=\mathbf{J}_{k}\boldsymbol{\Sigma}_{k}^{\prime}\mathbf{J}_{k}^{T}$, where $\mathbf{J}_{k}$ approximates the projective transformation at the Gaussian center. The view-dependent appearance is modeled using spherical harmonics and alpha composition.
Our Surgical Gaussian Surfels adapt the 3D Gaussian Splatting approach by using surface-aligned elliptical splats, optimized for rendering deformable surgical scenes. Inspired by [6], each surfel $\psi_{k}$ is parameterized by its position $p_{k}\in\mathbb{R}^{3}$, two orthonormal tangential vectors $t_{u},t_{v}\in\mathbb{R}^{3}$, and scaling factors $s_{u},s_{v}\in\mathbb{R}$ that control the size of the Gaussian. The inherent primitive normal is shown in Fig. 2 as $n_{k}=t_{u}\times t_{v}$ represents the direction of the steepest change in surfel density. This formulation enables direct gradient-based optimization, ensuring accurate surfel alignment with the surface of the surgical tissue.
The position of any point on the 2D Gaussian in world space is given by $P(u,v)=p_{k}+s_{u}t_{u}u+s_{v}t_{v}v,\ u,v\in\mathbb{R}$, where $(u,v)$ are local coordinates in the tangent plane. Using homogeneous coordinates, the transformation of the Gaussian from local to world space is represented as:
$H=\begin{bmatrix}s_{u}t_{u}&s_{v}t_{v}&0&p_{k}\\ 0&0&0&1\end{bmatrix},$ (1)
where $H\in\mathbb{R}^{4\times 4}$ encodes scaling, rotation, zeroed $\mathbf{z}$-component, and translation. As such, $\mathbf{R}_{k}\operatorname{diag}(\mathbf{s}_{k})=H_{k}[:,:3]$. It then follows that the density of each Surgical Gaussian Surfel primitive is modeled as:
$\mathcal{G}_{k}(u,v)=\exp\left(-\frac{u^{2}+v^{2}}{2}\right).$ (2)
Gaussian surfel deformation and rasterization. We use a lightweight MLP to model the deformation field of Gaussian Surfel primitives [17]. The MLP accepts the high-frequency positional and time-encoded [12] surfel primitives corresponding to the current frame as input. The deformation model learns adjustments for various Gaussian properties, such as position, scaling, and rotation, to capture the motion within the scene. Similar to [6], our Surgical Gaussian Surfels are rasterized following a depth-sorted volumetric alpha blending approach:
$c(x)=\sum_{k}c_{k}\alpha_{k}\hat{\mathcal{G}}_{k}(u(x))\prod_{j=1}^{k-1}\left(1-\alpha_{j}\hat{\mathcal{G}}_{j}(u(x))\right),$ (3)
where $c_{k}$ is the color of $k$-th Gaussian, $\alpha_{k}$ its opacity, and $\hat{\mathcal{G}}_{k}$ its adjusted Gaussian value. The process terminates once accumulated opacity reaches saturation.
3.2 PIMI Gaussian Surfel Primitive Initialization
Xie et al. [17] initialize a surgical scene point cloud by projecting visible pixels from depth maps and binary masks into 3D space. However, when depth data are incomplete, this results in noticeable gaps in the reconstructed point cloud and inaccuracies in the initialization due to color and geometric incongruencies.
To mitigate these issues, we propose an initialization method, termed Projection-Based Iterative Mask Integration (PIMI), which improves the quality and completeness of the initial point cloud by leveraging confidence-weighted depth aggregation and aggregated color mapping. The depth map $D_{t}$ for each frame is pre-processed to create a confidence-weighted depth map $D^{*}$ using a thresholded confidence mask, which ensures the exclusion of noisy depth data. Simultaneously, we aggregate the color maps $C_{t}$ to compute the refined color map $C^{*}$. This aggregation scheme at timestep $t$ is defined as:
$D^{*}=\frac{\sum_{t=1}^{T}D_{t}\cdot W_{t}}{\sum_{t=1}^{T}W_{t}},\quad C^{*}=\frac{\sum_{t=1}^{T}C_{t}\cdot W_{t}}{\sum_{t=1}^{T}W_{t}}$ (4)
where $W_{t}$ is the confidence mask derived by identifying reliable depth values within the 2nd to 99th percentile range of the depth distribution. Using $D^{*}$ and $C^{*}$, we generate the refined point cloud $P^{*}$. Each 2D pixel coordinate $(x,y)$ is projected into 3D space based on the intrinsic camera parameters $K$ as:
$p(x,y)=D^{*}(x,y)\cdot K^{-1}\begin{bmatrix}x\\ y\\ 1\end{bmatrix},\quad P^{*}=\{(p_{i},C_{i}^{*})\}_{i=1}^{N},$ (5)
where $P^{*}$ is the final point cloud, with $p_{i}$ as the 3D points and $C_{i}^{*}$ as their RGB colors encoded via spherical harmonics coefficients. This sequential integration of depth and color ensures geometric consistency and accurate color representation of the projected surgical scene.
3.3 Optimization
To capture fine details of surgical tissue and render areas occluded by surgical tools, we invert the binary mask in occluded areas and employ a weighted total loss consisting of photometric loss [14], total variation (TV) loss [11], deformation regularization, VGG16 [15] perceptual loss, and normal consistency:
$L=\lambda_{\text{1}}L_{\text{photo}}+\lambda_{\text{2}}L_{\text{tv}}+\lambda_{\text{3}}L_{\text{pos}}+\lambda_{\text{4}}L_{\text{cov}}+\lambda_{\text{5}}L_{\text{per}}+\lambda_{\text{6}}L_{\text{n}}.$ (6)
Deformation regularization. In single-view surgical video, surfel motion can become unstable without constraints. To ensure coherent yet flexible deformation, we use a deformation regularization that maintains relative distances between neighboring surfels before and after deformation. Given a surfel $\psi_{k}$, we enforce consistency across its $K$ nearest neighbors (where $K=5$[17] in both position and scaling:
$L_{\text{pos}}=\sum_{k}\sum_{j\in\mathcal{N}(k)}\left\|d(\mathbf{x}^{(k)},\mathbf{x}^{(j)})-d(\hat{\mathbf{x}}^{(k)},\hat{\mathbf{x}}^{(j)})\right\|_{1},$ (7)
$L_{\text{cov}}=\sum_{k}\sum_{j\in\mathcal{N}(k)}\left\|d(\mathbf{s}^{(k)},\mathbf{s}^{(j)})-d(\hat{\mathbf{s}}^{(k)},\hat{\mathbf{s}}^{(j)})\right\|_{1}.$ (8)
where $d(\cdot,\cdot)$ computes the Euclidean distance, and $\mathbf{x},\hat{\mathbf{x}}$ and $\mathbf{s},\hat{\mathbf{s}}$ denote initial and deformed surfel positions and scaling factors, respectively.
Normal consistency. The normal consistency term enforces alignment between the rendered normals and the implicit normals derived from the geometric properties of the scene. This regularization ensures that all Gaussian surfels are locally aligned with the tissue surfaces. The loss is defined as:
$L_{\text{n}}=\alpha\cdot(1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{expected}}))+\beta\cdot(1-\cos(\mathbf{n}_{\text{rendered}},\mathbf{n}_{\text{median}})),$ (9)
where $\mathbf{n}_{\text{rendered}}$, $\mathbf{n}_{\text{expected}}$, and $\mathbf{n}_{\text{median}}$ are the rendered normals, the expected normals from surface geometry (i.e., gradient of the depth map), and the normals of surfels at the median intersection point along the viewing ray, respectively. The weights $\alpha$ and $\beta$ balance the contributions of the two alignment targets.
Table 1: Quantitative comparisons of our method SGS with EndoNeRF [16], EndoSurf [20], LerPlane [18], EndoGaussian [10], and SurgicalGaussian (SG) [17].
Dataset Methods "pulling" "cutting" Rendering
LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ FPS↑ GPU↓
EndoNeRF EndoNeRF 0.080 29.088 0.930 0.106 26.792 0.901 0.04 15 GB
EndoSurf 0.111 36.919 0.961 0.103 35.551 0.955 0.05 17 GB
LerPlane 0.085 36.272 0.936 0.114 33.774 0.901 1.02 20 GB
EndoGaussian 0.089 36.429 0.951 0.102 35.773 0.956 190 2 GB
SG 0.049 38.783 0.970 0.062 37.505 0.961 80 4 GB
Ours 0.027 39.059 0.971 0.027 37.609 0.965 66 2 GB
"intestine" "liver" FPS↑ GPU↓
StereoMIS EndoNeRF 0.153 30.510 0.833 0.316 27.370 0.680 0.06 13 GB
EndoSurf 0.204 29.660 0.853 0.248 28.941 0.820 0.08 14 GB
LerPlane 0.206 29.441 0.822 0.254 28.852 0.793 1.45 19 GB
EndoGaussian 0.213 29.024 0.805 0.295 26.174 0.728 200 2 GB
SG 0.145 31.496 0.890 0.135 31.668 0.893 140 3 GB
Ours 0.048 34.669 0.936 0.047 34.390 0.934 81 2 GB
Table 2: Ablation study on EndoNeRF [16] dataset.
"pulling" "cutting"
[…]
w/o Perceptual loss 0.028 38.911 0.970 0.032 37.307 0.962
w/o Photometric loss 0.028 38.852 0.970 0.043 35.566 0.957
w GIDM initialization 0.027 38.907 0.970 0.031 37.380 0.963
Full model (PIMI init) 0.027 39.059 0.971 0.027 37.609 0.965
4 Experiments
Datasets and evaluation. We evaluate our method on two publicly available datasets, EndoNeRF [16] and StereoMIS [5]. The EndoNeRF dataset captures two cases of an in-vivo DaVinci robot prostatectomy from a single viewpoint. We evaluate on four video sequences of the StereoMIS dataset, which shows porcine subject breathing and tissue deformation during a liver and intestinal procedure. We estimate scene depth on the StereoMIS dataset using a pre-trained Omnidata [3] model and tool occlusion binary masks using a pre-trained SAM2 [13] model. We measure the surgical scene reconstruction capabilities of SGS using three standard metrics: LPIPS, PSNR, and SSIM. Our method achieves state-of-the-art (SOTA) results on all metrics.
Implementation details. All experiments are carried out on an NVIDIA RTX A6000, using the PyTorch framework. Adam [9] optimizer is used during training. The loss weights in Eq. 6 were set empirically to: $\lambda_{\text{1}}=0.7,\ \lambda_{\text{2}}=0.006,\ \lambda_{\text{3}}=1.0,\ \lambda_{\text{4}}=200.0,\ \lambda_{\text{5}}=1.0,\ \lambda_{\text{6}}=0.0001.$
EndoNeRF EndoGaussian LerPlane SG Ours Reference
Figure 3: Qualitative comparisons of our method (SGS) with EndoNeRF [16], EndoSurf [20], LerPlane [18], EndoGaussian [10], and SG [17].
4.1 Main Results and Ablation
We compare SGS against existing methods for real-time rendering of surgical scenes in the EndoNeRF [16] and StereoMIS [5] datasets. SDF-based approaches such as EndoSurf [20] require separate MLPs for deformation, geometry, and appearance, which increases computational complexity and limits real-time performance. Neural LerPlane [18] factorizes 4D scenes into 2D planes to reduce overhead, but struggles to capture complex deformations. EndoGaussian [10] shifts towards explicit depth-aware point clouds but relies heavily on pre-initialized geometry, making it less adaptable to dynamic changes. Lastly, SurgicalGaussian [17] introduces a deformation network that decouples motion and geometry, but struggles to render fine details in the surgical scene. SGS outperforms these previous methods in all datasets, achieving the lowest LPIPS, the highest PSNR and the highest SSIM, while qualitatively capturing specularity and anatomical features more accurately, as shown in Tab. 1 and Fig. 3.
Ablation study. We perform ablation studies in a "leave-one-out" fashion on the EndoNeRF [20] dataset. To assess the impact of our PIMI Gaussian initialization, we replace it with the prior SOTA GIDM [17] point cloud initialization. We also provide test-time results without perceptual, photometric, and TV loss. The results are summarized in Tab. 2. Homodirectional view-space positional gradients [19] are used during Gaussian Surfel densification for all experiments.
5 Conclusion
In this article, we introduce Surgical Gaussian Surfels (SGS), an explicit rendering method that takes advantage of the accurate geometry reconstruction capabilities of surface-aligned anisotropic splats. We utilize scale-constrained elliptical primitives to produce high-fidelity renders of surgical scenes. We define surface normals as the direction of the steepest density change, thereby generating accurate surgical tissue surface normals without normal priors. By leveraging both Projection-Based Iterative Mask Integration (PIMI) for robust point-cloud initialization and Gaussian Surfel densification based on homodirectional view-space positional gradients, our method achieves state-of-the-art reconstruction quality with real-time rendering performance. SGS has the potential to advance surgical scene understanding and support a range of intraoperative applications, including surgical navigation and robot-aided surgical procedures.